# Exploiting the Final Component of Generator Architectures for AI-Generated Image Detection

### v1

@@ 3.2.1 Construction from the three representative components
To investigate the generalization of final component $\varphi$ in generated trace detection (Eq. 1), we identify three representative final components from open-source generators in our taxonomy as shown in Fig.1: Stable Diffusion 2.1 (VAE.decoder-1.1:Latent diffusion), JanusPro (VQ.de-tokenizer-2.2:Next-token auto-regression), and PixelFlow (Diffusion.denoiser-3.3:End-to-end super-resolution). For brevity, we refer to these specific components as VAE_1.1, VQ_2.2, and Super-resolution_3.3 respectively in the following sections. As discussed in Property 1 (Sec. 3), the encoder $\mathcal{E}^{*}$, which is pretrained jointly with $\varphi^{*}$, is used to approximate the latent input $\mathbf{z}$.
Algorithm 1 presents the pseudocode for constructing $\varphi^{*}(\mathcal{E}^{*}(\mathbf{x}))$ using the VAE_1.1. The implementation requires only a few lines of code and is significantly more efficient than executing the full generative pipeline. For VQ_2.2, the de-tokenizer acts as $\varphi^{*}$ and the image tokenizer as $\mathcal{E}^{*}$ to approximate image tokens $\mathbf{z}$, as outlined in Algorithm 2. In Super-resolution_3.3, the input $\mathbf{z}$ is a noisy version of the real image, $\alpha\mathbf{x}+\beta\epsilon_{t}$, where $\epsilon\sim\mathcal{N}(0,I)$. The noise level can be adjusted by selecting the starting timestep $t$ for denoising (Algorithm 3).
Figure 3: Performance comparison of our model trained on the final component of SD2
Figure 4: Performance comparison of our model trained on the final component of HiDream
Figure 5: Feature visualization of the three final components.
@@ 3.2.2 Sparse Sampling from Combined Constructions
The three final components differ substantially in both functionality and implementation. Consequently, the constructed samples $\varphi^{*}(\mathcal{E}^{*}(\mathbf{x}))$ from each reflect distinct generative traces. We further investigate whether combining a small number of representative fake samples from each component can better capture the overall trace space and improve generalization. Algorithm 4 outlines the procedure, applying K-Medoids clustering selects $k$ representative samples from the constructed outputs of each component, as described in Section 3.2.1. The selection is performed in a reduced-dimensional feature space. Specifically, the generated samples are first encoded using a pretrained DINOv3 backbone, from which 2,048-dimensional feature vectors are extracted from the final layer. K-Medoids selects $k$ samples such that most other samples are close to one of them in Euclidean distance. The selected subsets from all three components are combined to form the final training set of $\varphi^{*}(\mathcal{E}^{*}(\mathbf{x}))$ for detector training.
@@ 3.3 Detection model — $\theta$
Algorithm 4 Negative sample selection
Input: Constructed image sets $\mathcal{D}_{1}\cup\mathcal{D}_{2}\cup\mathcal{D}_{3}=\{\hat{\mathbf{x}}\}_{VAE}\cup$ $\{\hat{\mathbf{x}}\}_{VQ}\cup\{\hat{\mathbf{x}}\}_{sup}$, the number of samples to be selected $k$
Output: The sparse constructed image set $\{\hat{\mathbf{x}}\}$
1: dinov3 = torch.hub.load(“dinov3_vith16plus”)
2:for $i=1$ to 33 do
4:  dist = sklearn.metrics.euclidean_distances(features)
5:  $\mathcal{M}_{i}$= sklearn.kmedoids.fasterpam(dist, $k$) //a set of $k$ samples of $\hat{\mathbf{x}}$
6:end for
7:return $\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{3}$
Based on Property 2 discussed in Section 3, instead of loading $\mathbf{x}$ and computing $\varphi^{*}(\mathcal{E}^{*}(\mathbf{x}))$ at each training step, we directly input the real image set (positive samples) $\{\mathbf{x}\}$ and the constructed image set $\{\hat{\mathbf{x}}\}$ generated by Algorithms 14, with standard shuffling applied to both sets during SGD training. Since the detector $\theta$ aims to distinguish subtle differences between real and reconstructed images, and object detection tasks inherently emphasize multi-scale spatial features, we adopt DINOv3—a foundation model pretrained for object detection—as the backbone to better capture fine-grained artifacts. A single fully connected layer is appended to the backbone, and all layers are fine-tuned during training. As our focus is on analyzing the effect of the final generative component, we employ a pretrained visual backbone with minimal modifications rather than designing a complex detector.
@@ 4.1 Model implementation & baselines
The real image set used for constructing the training dataset is MS-COCO 2014 [21], which contains 118,000 images depicting diverse indoor and outdoor scenes with resolutions ranging from 384 to 880 pixels in width (height). The constructed training and corresponding test sets are described separately for each of the following experimental questions, with all pretrained models referenced in the Appendix.
@@
For our AI-generated image detector, we use the pretrained backbone of DINOv3111https://huggingface.co/facebook/dinov3-vith16plus-pretrain-lvd1689m with an additional fully connected layer for binary classification. All layers are fine-tuned using standard cross-entropy loss with a learning rate of 5e-7. This implementation serves as the detector for all subsequent analyses.
@@
We compare our method against eight baseline approaches: AIDE [28] (ICLR 2025), C2P-CLIP [44] (AAAI 2025), CoDE [19] (ECCV 2024), RINE [23] (ECCV 2024), FatFormer [22] (CVPR 2024), NPR [18] (CVPR 2024), LGrad [10] (CVPR 2023), and DIRE [16] (ICCV 2023). As none of these methods report results on our testing sets, we re-evaluate all of them using their publicly released pretrained models and default hyperparameters. The source code links, parameters, and models used are detailed in the Appendix. Performance is assessed using two standard metrics: classification accuracy and average precision (AP).
4.2 Results & discussions
Table 2.Detection results across different category of generators.
vae.decoder based diffusion.denoiser.super-resolution based vq.decoder based
SD1.3 SD1.4 SD XL Flux SD 3.5 DALL·E 2 DALL·E 3 Glide Emu3 LlamaGen
Acc AP Acc AP Acc AP Acc AP Acc AP Acc AP Acc AP Acc AP Acc AP Acc AP
DIRE (ICCV2023) 48.75 0.43 48.65 0.43 50.00 0.59 61.66 0.73 55.65 0.59 56.45 0.67 50.35 0.54 49.50 0.49 53.65 0.57 54.05 0.59
LGrad (CVPR2023) 54.45 0.46 55.85 0.48 52.60 0.48 81.91 0.92 77.29 0.87 55.65 0.56 47.65 0.40 47.45 0.52 54.80 0.48 60.70 0.63
NPR (CVPR2024) 50.70 0.51 50.60 0.52 52.45 0.56 90.00 0.92 88.04 0.90 49.25 0.50 50.35 0.51 48.80 0.51 51.70 0.53 52.25 0.55
FatFormer (CVPR2024) 52.97 0.54 53.72 0.54 66.67 0.75 55.32 0.90 75.86 0.95 54.12 0.56 35.88 0.34 67.77 0.76 63.72 0.66 77.21 0.87
RINE (ECCV2024) 87.60 0.96 87.15 0.96 74.60 0.86 51.70 0.87 96.49 0.99 85.90 0.94 42.35 0.31 63.95 0.72 62.10 0.70 90.65 0.98
CoDE (ECCV2024) 97.75 0.97 97.30 0.97 75.20 0.75 71.09 0.72 76.35 0.75 54.60 0.54 73.15 0.73 80.75 0.80 77.45 0.77 90.50 0.90
C2P-CLIP (AAAI2025) 75.15 0.93 76.70 0.93 77.70 0.95 53.96 0.97 80.88 0.99 55.60 0.77 63.20 0.90 86.65 0.97 68.70 0.91 83.80 0.98
AIDE (ICLR2025) 57.60 0.60 56.70 0.60 62.70 0.64 98.50 0.99 96.45 0.99 59.75 0.43 14.15 0.31 59.75 0.77 62.50 0.57 63.25 0.85
Ours_vae_1.1 99.70 1.0 99.70 1.0 99.70 1.0 99.17 .9998 99.96 1.0 99.60 1.0 99.70 .9999 98.85 .9992 99.70 .9999 99.70 .9999
Ours_vq_2.2 99.85 1.0 99.85 1.0 99.85 1.0 99.19 .9997 99.82 .9999 99.80 1.0 99.85 .9999 98.75 .9984 99.85 .9999 99.85 1.0
Ours_super-res_3.3 99.90 1.0 99.80 1.0 99.90 1.0 99.89 1.0 99.86 1.0 99.35 .9991 99.85 1.0 98.60 .9985 99.90 1.0 99.90 1.0
Ours_sparse 99.15 .9990 99.15 .9990 99.15 .9988 98.95 .9828 98.98 .9958 99.80 .9987 99.15 .9991 98.05 .9947 99.15 .9990 99.15 .9982
In experiments, we aim to address the following questions:
Q1: Are traces from the final component alone sufficient to detect images generated by its full pipeline?
To verify this, we train two detectors on samples from the VAE decoders of SD2 and HiDream-I1 and evaluate them on images produced by their respective full generation pipelines.
Dataset: (1) SD2. The training samples comprise real MS-COCO images and corresponding fake images constructed via Algorithm 1 using Stable Diffusion 2.1. The test set includes 1,000 real images and 1,000 SD2-generated images from Synthbuster [^2] https://zenodo.org/records/10066460. (2) HiDream-I1. The fake training samples are constructed in same procedure but using the HiDream-I1 model, and the test set is generated from the same 1,000 prompts as Synthbuster.
Fig.s 5 and 5 show that both detectors perform strongly. For SD2, baselines vary widely. CoDE and C2P-CLIP achieve relatively high results because their training sets include SD2-generated images, making this an in-distribution evaluation. In contrast, HiDream-I1 evaluation is fully zero-shot. For the AP metric, our detector achieves nearly perfect scores, with logits for real samples approaching 1 and those for fake samples approaching 0. These results demonstrate that training with the generator’s final component generalizes effectively to the generator’s full outputs.
Q2: Are traces from the final component sufficient to detect images from full pipelines across categories?
Table 3: Detection results on unknown generators.
Unknown Mixture Wild
@@ 4.2 Results & discussions
Acc AP Acc AP Acc AP Acc AP Acc AP Acc AP
DIRE 50.8 0.55 53.9 0.61 44.38 0.42 57.19 0.62 56.56 0.66 47.74 0.79
LGrad 47.25 0.42 57.7 0.53 44.01 0.69 57.80 0.65 66.88 0.77 40.86 0.61
NPR 47.30 0.47 52.35 0.55 52.37 0.58 63.80 0.67 74.06 0.84 48.46 0.68
FatFormer 60.12 0.62 44.43 0.42 71.2 0.81 69.47 0.81 54.83 0.75 42.79 0.77
RINE 92.10 0.99 55.45 0.59 74.77 0.97 71.20 0.81 51.56 0.55 55.93 0.62
CoDE 60.65 0.60 76.90 0.77 74.75 0.74 65.8 0.61 72.19 0.67 68.41 0.78
C2P-CLIP 59.45 0.82 52.85 0.69 74.77 0.91 68.40 0.79 54.37 0.71 47.27 0.75
AIDE 14.40 0.32 53.45 0.40 74.78 0.87 70.87 0.83 50.00 0.56 52.78 0.59
Ours_vae_1.1 99.70 1.0 99.70 .9999 98.83 .9997 96.87 .9968 97.81 .9899 98.34 .9993
Ours_vq_2.2 99.70 .9999 99.45 .9996 98.10 .9982 95.40 .9980 95.31 .9928 98.34 .9986
Ours_sup-res 98.50 .9996 99.70 1.0 94.38 .9978 98.67 .9964 97.19 .9946 97.39 .9998
Ours_sparse 98.00 .9973 99.10 .9990 97.90 .9991 97.60 .9951 98.44 .9944 99.31 1.0
@@
amuse- park car dealership electric station stadium
Acc AP Acc AP Acc AP Acc AP
50.49 0.86 50.19 0.84 50.40 0.84 50.58 0.94
48.85 0.36 50.00 0.42 50.15 0.47 50.02 0.43
46.89 0.43 65.52 0.84 65.65 0.93 56.16 0.85
51.68 0.70 50.88 0.83 52.81 0.79 51.66 0.84
69.07 0.83 51.97 0.50 64.19 0.74 68.16 0.78
86.66 0.82 88.40 0.84 84.50 0.80 78.37 0.73
58.00 0.76 76.60 0.96 71.71 0.88 58.74 0.94
52.90 0.58 61.71 0.85 61.11 0.76 52.32 0.72
99.33 1.0 96.43 .9996 96.08 .9987 96.50 .9997
96.77 1.0 87.52 .9998 90.99 .9989 98.89 .9999
99.67 .9997 97.94 .9962 96.50 .9917 97.34 .9943
99.61 .9999 98.91 .9994 98.55 .9982 99.59 .9998
We train detectors on the three representative final components (Section 3.2.1) and an additional detector on the sparse image set (Section 3.2.2), then evaluate them on images from different categories in our taxonomy.
Dataset: Training sets. For the VAE_1.1, VQ_2.2, and Super-Resolution_3.3 components, real images are from MS-COCO, and fake samples are constructed following Algorithms 13. These three types of final components exhibit distinct generation traces, visualized in Fig. 5. Features are extracted from the third layer of the pretrained DINOv3 backbone and project into three dimensions. The resulting distributions appear mixed yet remain visually separable. The sparse training set is selected using Algorithm 4, with $k=100$. As shown in Fig. 6, the selected data points (highlighted in red) span the entire space. Testing set. The testing set consists of image subsets from the Synthbuster benchmark, generated by SD1.3, SD1.4, SD XL, DALL·E 2, DALL·E 3, and Glide. We categorize these subsets according to our taxonomy, as summarized in Table 6. For the VQ decoder category, we generate fake images from two autoregressive generators, Emu3 and LlamGen, using the 1,000 Synthbuster prompts. For emerging generators Flux-dev and SD3.5, we produce 10,000 fake images each. As paired real images are unavailable, we use 9,600 real images from the PASCAL VOC training split [11] for evaluation.
Figure 6: Sparse sample selection.
Table 6 reports detection results across generator categories. Each row in the last group corresponds to a detector trained on one of the four training sets—VAE_1.1, VQ_2.2, Super-resolution_3.3, or the sparse set. Our approach performs strongly within the same category and generalizes well across categories. For example, on DALL·E 2—where most baselines are near random (e.g., NPR: 49.20%, LGrad: 55.65%, FatFormer: 54.12%)—only CLIP-based models like RINE (85.90%) and AIDE (59.75%) show moderate gains, benefiting from DALL·E 2’s CLIP-based design. Despite not relying on CLIP features, our method achieves substantially higher accuracy, highlighting robust generalization to unseen generator architectures.
Notably, the detector trained on the sparse set, using only 100 fake samples from each of the three components (300 fake and 300 real images in total), achieves performance comparable to our three models trained on individual components with the full MS-COCO dataset. This small training set provides further evidence of the generalization capability of our approach and offers a practical advantage, as generating large-scale training images is costly.
Q3: Are traces from the final component sufficient to detect images from wild or unknown generators?
The detectors trained for Q2 are directly evaluated on testing sets for this question.
Dataset: Testing set. (1) Unknown generators — Firefly and Midjourney test sets from Synthbuster. Both are commercial generators lacking open-source code or public documentation of their architectures or components. (2) Mixed generators—FakeBench  [^3] https://github.com/Yixuan423/FakeBench benchmark contains 3,000 real images and 3,000 fake images generated by multiple models. (3) In-the-wild generation—WildRF  [^4] https://vision.huji.ac.il/ladeda/ comprises three subsets: Reddit, Facebook, and Twitter, each containing real and synthetic images sourced from social media platforms with unknown origins.
Table 4 reports the results. Baseline methods exhibit highly inconsistent results across benchmarks. For instance, RINE performs well on Firefly (92.10% accuracy, 0.99 AP) but drops sharply on Midjourney (55.45%) and Facebook (51.56%), while AIDE misclassifies most samples—achieving only 14.40% accuracy. This instability highlights the challenge of real-world generalization, particularly in true zero-shot scenarios where no prior information about the generators is available. In contrast, our detector consistently surpasses all baselines and maintains robust, stable performance across all subsets, demonstrating that leveraging traces from the final component yields strong generalization to diverse architectures and challenging real-world conditions.
Q4: Does detection performance on a full generator remain valid after fine-tuning the generator?
The detectors trained for Q2 are directly evaluated on testing sets for this question.
Dataset: Testing set — SatelliteDiffusion[15] dataset is generated using a domain-specific model fine-tuned from SD2, conditioned on satellite metadata. Real images are sourced from high-resolution satellite imagery, while synthetic counterparts are produced using textual descriptions of the real samples. Four subsets: amusement park, car dealership, electric substation, and stadium are tested, each including around 2,000 real and 2,000 generated images.
Table 4 presents the comparison. Our models attain results closely matching those reported for the original SD2 evaluations (Fig. 5, Table 6), demonstrating that detection remains effective for generators fine-tuned on domain-specific images and conditions.

### v2

@@ 4.1.1 Construction from the three representative components
In our taxonomy (Fig. 1), the three primary categories—VAE.decoder, VQ.de-tokenizer, and Diffusion.denoiser—are defined based on the implementation of the final component, while subcategories are further determined by the input space and functional context in which the component operates. Consequently, all components within the same primary category can share the same algorithmic pipeline for sample construction.
For brevity, we refer to components in these categories as VAE, VQ, and Diffusion respectively in the following sections. As discussed in Property 1, the encoder $\mathcal{E}^{*}$, which is pretrained jointly with $\varphi^{*}$, is used to approximate the latent input $\mathbf{z}$. Algorithm 1 presents the pseudocode for constructing $\varphi^{*}(\mathcal{E}^{*}(\mathbf{x}))$ using the VAE. The implementation requires only a few lines of code and is significantly more efficient than executing the full generative pipeline. For VQ, the de-tokenizer acts as $\varphi^{*}$ and the image tokenizer as $\mathcal{E}^{*}$ to approximate image tokens $\mathbf{z}$, as outlined in Algorithm 2. In Diffusion, the input $\mathbf{z}$ is a noisy version of the input, $\alpha\mathbf{x}+\beta\epsilon_{t}$, where $\epsilon\sim\mathcal{N}(0,I)$. The noise level can be adjusted by selecting the starting timestep $t$ for denoising (Algorithm 3).
@@ 4.1.2 Sparse sampling from combined constructions
As the final components of the three categories differ substantially in both functionality and implementation, the constructed samples $\varphi^{*}(\mathcal{E}^{*}(\mathbf{x}))$ from each may exhibit distinct generative traces. We further investigate whether combining a small set of representative fake samples from each component can capture the overall trace space and achieve effective generalization. Algorithm 4 outlines the procedure, applying K-Medoids clustering selects $k$ representative samples from the constructed outputs of each component. The selection is performed in a reduced-dimensional feature space. Specifically, the generated samples are first encoded using a pretrained DINOv3 backbone [39], from which 2,048-dimensional feature vectors are extracted from the final layer. K-Medoids selects $k$ samples such that most other samples are close to one of them in Euclidean distance. The selected subsets from all three components are combined to form the final training set of $\varphi^{*}(\mathcal{E}^{*}(\mathbf{x}))$ for detector training.
@@ 4.2 Detection model — $\theta$
Based on the Property 2, instead of loading $\mathbf{x}$ and computing $\varphi^{*}(\mathcal{E}^{*}(\mathbf{x}))$ at each training step, we directly input the real image set $\{\mathbf{x}\}$ and the constructed image set $\{\hat{\mathbf{x}}\}$ generated by Algorithms 14, with standard shuffling applied to both sets during SGD training. Since the detector $\theta$ aims to distinguish subtle differences between real and reconstructed images, and object detection tasks inherently emphasize multi-scale spatial features, we adopt DINOv3—a foundation model pretrained for object detection—as the backbone to better capture fine-grained artifacts. A single fully connected layer is appended to the backbone, and all layers are fine-tuned during training. As our focus is on analyzing the effect of the final generative component, we employ a pretrained visual backbone with minimal modifications rather than designing a complex detector.
Table 1. Detection results across different category of generators.
diffusion.denoiser based vq.de-tokenizer based vae.decoder
DALL·E 2 DALL·E 3 Glide Emu3 LlamaGen HiDream
Acc% AP% Acc% AP% Acc% AP% Acc% AP% Acc% AP% Acc% AP%
DIRE 56.45 67.05 50.35 53.53 49.50 49.27 53.65 56.73 54.05 59.26 51.40 51.21
LGrad 55.65 55.73 47.65 40.41 47.45 51.83 54.80 48.03 60.70 63.08 40.05 39.06
NPR 49.25 50.18 50.35 50.48 48.80 51.43 51.70 53.05 52.25 55.48 47.75 45.51
FatFormer 54.12 56.02 35.88 33.95 67.77 75.75 63.72 66.10 77.21 87.73 34.48 31.28
RINE 85.90 94.05 42.35 30.83 63.95 72.09 62.10 69.68 90.65 98.36 44.15 39.20
CoDE 54.60 54.26 73.15 72.76 80.75 80.36 77.45 77.06 90.50 90.11 71.45 71.06
C2P-CLIP 55.60 76.87 63.20 90.45 86.65 96.53 68.70 90.60 83.80 97.85 51.20 51.02
AIDE 59.75 42.26 14.15 31.34 59.75 76.71 62.50 56.86 63.25 84.55 48.30 37.95
BFree 93.30 99.16 96.22 99.64 68.61 93.81 99.05 99.97 99.15 99.98 81.40 93.19
Ours_VAE 99.60 100.0 99.70 99.99 98.85 99.92 99.70 99.99 99.70 99.99 99.15 99.94
Ours_VQ 99.80 100.0 99.85 99.99 98.75 99.84 99.85 99.99 99.85 100.0 97.15 99.51
Ours_Diffusion 99.35 99.91 99.85 100.0 98.60 99.85 99.90 100.0 99.90 100.0 88.95 96.95
Ours_Sparse 99.80 99.87 99.15 99.91 98.05 99.47 99.15 99.90 99.15 99.82 96.80 99.21
vae.decoder based
SD1.3 SD1.4 SD XL SD2 Flux SD 3.5
Acc% AP% Acc% AP% Acc% AP% Acc% AP% Acc% AP% Acc% AP%
DIRE 48.75 43.24 48.65 43.38 50.00 58.90 47.45 40.37 61.66 73.10 55.65 58.00
LGrad 54.45 46.46 55.85 48.14 58.65 53.26 48.75 42.80 81.91 91.81 77.29 87.08
NPR 50.70 51.05 50.60 51.69 52.45 55.54 43.50 46.23 90.00 92.34 88.04 89.82
FatFormer 52.97 53.60 53.72 54.07 66.67 75.43 51.57 54.41 55.32 89.83 75.86 95.39
RINE 87.60 95.96 87.15 95.95 74.60 85.56 70.30 80.74 51.70 86.61 96.49 99.93
CoDE 97.75 97.36 97.30 96.91 75.20 74.81 88.05 87.66 71.09 72.17 76.35 75.26
C2P-CLIP 75.15 92.93 76.70 93.03 77.70 94.67 69.15 92.30 53.96 94.58 80.88 98.87
AIDE 57.60 60.25 56.70 60.14 62.70 63.98 53.45 42.09 98.50 99.80 96.45 99.72
BFree 99.25 99.99 99.15 99.99 99.25 99.98 99.80 99.93 69.62 93.62 98.74 99.94
Ours_VAE 99.70 100.0 99.70 100.0 99.70 100.0 99.70 100.0 99.17 99.98 99.96 100.0
Ours_VQ 99.85 100.0 99.85 100.0 99.85 100.0 99.40 99.90 99.19 99.97 99.82 99.99
Ours_Diffusion 99.90 100.0 99.80 100.0 99.90 100.0 99.85 100.0 99.89 100.0 99.86 100.0
Ours_Sparse 99.15 99.90 99.15 99.90 99.15 99.88 99.10 99.80 98.95 98.28 98.98 99.58
@@ 5.1 Model implementation & baselines
Training dataset The real image set used to construct the training dataset in this study is MS-COCO 2014 [28], which contains 118,000 images depicting diverse indoor and outdoor scenes, with resolutions ranging from 384 to 880 pixels in width (or height).
To validate the generalizability of the generator’s final component, we select one representative final component from open-source generators for each of the three categories in Fig. 1: VAE.decoder, VQ.detokenizer, and Diffusion.denoiser. Specifically, Stable Diffusion 2.1 represents the VAE.decoder category and is used to construct fake samples following Algorithm 1; JanusPro represents the VQ.detokenizer category and generates samples according to Algorithm 2; and PixelFlow represents the Diffusion.denoiser category and follows Algorithm 3. All pretrained encoders $\mathcal{E}^{*}$ and decoders $\varphi^{*}$ are detailed in the Appendix. For brevity, we refer to the corresponding training sets as VAE, VQ, and Diffusion in the following discussion.
These three types of final components exhibit distinct generation traces, as visualized in Fig. 4. Features are extracted from the third layer of a pretrained DINOv3 backbone [39] and projected into three dimensions. Although the resulting distributions exhibit partial overlap, they remain visually separable. To further examine whether a small number of representative samples from each component can capture the overall trace space and achieve effective generalization, we construct an additional compact training set. Specifically, we select 100 samples from each of the VAE, VQ, Diffusion sets using Algorithm 4 with $k=100$, and refer to it as Sparse set in experiments. Fig. 4 visualizes the selection results, where the chosen samples (highlighted in red) are distributed throughout the entire feature space.
@@
Detector implementation To implement the AI-generated image detector, we adopt a pretrained DINOv3 backbone [39] with an additional fully connected layer for binary classification. All layers are fine-tuned using the standard cross-entropy loss in Eq. 1 with a learning rate of $5\times 10^{-7}$. We train separate detectors on the training sets VAE, VQ, Diffusion, and Sparse, and evaluate them on images from different categories in our taxonomy to analyze the effectiveness of the proposed method.
@@
Baselines We compare our method against nine baseline approaches: BFree [17] (CVPR 2025), AIDE [35] (ICLR 2025), C2P-CLIP [51] (AAAI 2025), CoDE [25] (ECCV 2024), RINE [30] (ECCV 2024), FatFormer [29] (CVPR 2024), NPR [24] (CVPR 2024), LGrad [14] (CVPR 2023), and DIRE [22] (ICCV 2023). As these methods either did not report results on all of our test sets or used different evaluation metrics, we re-evaluate each method using its publicly released pretrained model and default hyperparameters. Notably, BFree also evaluated these baselines on a subset of our test sets; our results and conclusions regarding the baseline methods are consistent with those reported in BFree. The source code links, configurations, and pretrained models used are provided in the Appendix. Performance is assessed using two standard metrics: classification accuracy (Acc) and average precision (AP).
Table 2: Detection results on wild generations.
Unknown architecture Mixture Wild generation
@@
Acc% AP% Acc% AP% Acc% AP% Acc% AP% Acc% AP% Acc% AP%
DIRE 50.80 55.06 53.90 60.56 44.38 41.72 57.19 61.93 56.56 66.02 47.74 78.72
LGrad 47.25 42.41 57.70 53.09 44.01 69.44 57.80 65.31 66.88 77.26 40.86 61.63
NPR 47.30 46.96 52.35 55.22 52.37 57.85 63.80 66.87 74.06 83.68 48.46 67.53
FatFormer 60.12 62.24 44.43 41.82 71.20 80.93 69.47 80.83 54.83 75.27 42.79 77.41
RINE 92.10 99.71 55.45 59.38 74.77 97.22 71.20 81.28 51.56 55.44 55.93 62.27
CoDE 60.65 60.28 76.90 76.51 74.75 73.94 65.80 61.12 72.19 67.50 68.41 78.34
C2P-CLIP 59.45 82.11 52.85 69.37 74.77 91.18 68.40 78.85 54.37 71.39 47.27 75.07
AIDE 14.40 31.74 53.45 40.43 74.78 86.76 70.87 82.52 50.00 55.96 52.78 59.40
BFree 98.85 99.91 98.70 99.93 80.81 100.0 85.20 94.94 95.31 98.98 95.75 99.42
Ours_VAE 99.70 100.0 99.70 99.99 98.83 99.97 96.87 996.8 97.81 98.99 98.34 99.93
Ours_VQ 99.70 99.99 99.45 99.96 98.10 99.82 95.40 99.80 95.31 99.28 98.34 99.86
Ours_Diffusion 98.50 99.96 99.70 100.0 94.38 99.78 98.67 99.64 97.19 99.46 97.39 99.98
Ours_Sparse 98.00 99.73 99.10 99.90 97.90 99.9 97.60 99.51 98.44 99.44 99.31 100.0
@@
amusement_park car_dealership electric_station stadium
Acc% AP% Acc% AP% Acc% AP% Acc% AP%
DIRE 50.49 85.96 50.19 84.49 50.40 83.85 50.58 94.11
LGrad 48.85 35.98 50.00 42.38 50.15 47.30 50.02 43.46
NPR 46.89 42.62 65.65 92.65 62.52 83.55 56.16 84.64
FatFormer 51.68 69.84 50.88 83.10 52.81 78.68 51.66 83.76
RINE 69.07 82.59 51.97 50.15 64.19 73.99 68.16 78.23
CoDE 86.66 82.00 88.40 83.67 84.50 79.53 78.37 73.16
C2P-CLIP 58.00 75.82 76.60 96.21 71.71 87.63 58.74 93.90
AIDE 52.90 57.74 61.71 84.69 61.11 75.62 52.32 72.37
BFree 95.48 100.0 95.57 99.97 90.75 99.95 88.95 99.98
Ours_VAE 99.33 100.0 96.43 99.96 96.08 99.87 96.50 99.97
Ours_VQ 96.77 100.0 87.52 99.98 90.99 99.89 98.89 99.99
Ours_Diffusion 99.67 99.97 97.94 99.62 96.50 99.17 97.34 99.43
Ours_Sparse 99.61 99.99 98.91 9.994 98.55 99.82 99.59 99.98
5.2 Results across different category of generators
Testing set (1) Synthbuster [3] is a challenging benchmark comprising 1,000 high-resolution real images from RAISE-1K [11] and 1,000 fake images generated by each of nine models: SD1.3, SD1.4, SD2, SDXL, DALL·E2, DALL·E3, Glide, Adobe Firefly and MidJourney v5, using the corresponding 1000 prompts. Based on our proposed taxonomy, these subsets are classified into the VAE.decoder and Diffusion.denoiser categories, as summarized in Table 1. As the architectures of Firefly and MidJourney are not disclosed, we treat them as wild generators and evaluate them in Section 5.3.
Since the release of Synthbuster, additional generators have emerged that are not included in the benchmark, including VQ.decoder-based generators and new VAE.decoder-based generators such as HiDream, Flux-dev, and SD3.5. To incorporate these, we generate fake images using two autoregressive VQ.decoder generators, Emu3 and LlamGen, as well as HiDream, all based on the same 1,000 Synthbuster prompts.
(2) Although the AI-generated images in these sub-testing sets come from various generators, the real image set is the same across all subsets. In other words, the model’s classification performance on the real images affects the results for all subsets when computing balanced binary classification accuracy. Moreover, 1,000 images for both the real and fake sets constitute a limited sample size and offer limited diversity. To evaluate more diverse semantic content and larger sample sizes, for Flux-dev and SD3.5, we use 9,600 images from the PASCAL VOC training split [15] as the real image set. For the fake image set, we generate 10,000 images from each generator using short descriptions produced by an LLM. Since the detector identifies generator traces rather than comparing real–AI image pairs, the semantic content differences between PASCAL VOC and the generated images do not affect the evaluation. Instead, the ability of the detector to correctly classify real images containing many small objects, as in PASCAL VOC, is a key aspect of assessing its robustness.